OpenStudy (anonymous):

If 5:m as 12:60, then m is equal to ______

Directrix (directrix):

If 5:m as 12:60, then as a proportion, 5/m = 12/60. 5/m = 1/5. Using the means extremes property, m = 5*5 = 25.

OpenStudy (anonymous):

yep